OpenStudy (anonymous):

help please 7^2 x (6-3)

OpenStudy (anonymous):

solving or simplifying?

OpenStudy (anonymous):

@nicole-0109

OpenStudy (anonymous):

I think's solving

OpenStudy (anonymous):

OK In that case do you know PEMDAS?

OpenStudy (anonymous):

Or Bedmas?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

It is the order in which to solve something. OK it stands for Please Excuse My Dear Aunt Sally Taking the first letter of each word gives you Parenthesis Exponents Multiplication & Dividing Adding and Subtracting

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So taking \(\huge 7^{2} x (6-3)\) Which do you solve first?

OpenStudy (anonymous):

The X is a times right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

OK In the future use the astracic * (type shift and 8) x can be used as a varible in algebra

OpenStudy (anonymous):

now taking \(\huge 7^{2}*(6-3)\) which do you solve first?

OpenStudy (anonymous):

Using pedmas?

OpenStudy (anonymous):

7^2 ??

OpenStudy (anonymous):

im wrong is (6-3)

OpenStudy (anonymous):

Close, rememeber Pedmas the first letter of everything is what you should solve first in this equation you can solve the parenthesis or exponent first, but in the future remember to solve parenthesesis firist but Ok solve for \( 7^{2}\)

OpenStudy (anonymous):

EXACTLLY! 6-3! sorry i was yping didn't see your reply

OpenStudy (anonymous):

that's ok

OpenStudy (anonymous):

OK Solve for 6-3

OpenStudy (anonymous):

Then solve for \(\huge7^{2}\) Do you know how to solve exponets?

OpenStudy (anonymous):

can u give me a example please

OpenStudy (anonymous):

Example of what?

OpenStudy (anonymous):

7^2=49 ??

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

So given \(\huge 49*(6-3)\) 6-3=?

OpenStudy (anonymous):

6-3=3

OpenStudy (anonymous):

Yup so that leaves us with \(\huge 49*3\)

OpenStudy (anonymous):

147

OpenStudy (anonymous):

Yup :) Congrats! \(\huge \color {royalblue} {Great!}\)

OpenStudy (anonymous):

Don't forget to best answer :)

OpenStudy (anonymous):

lol ok thank you could you help with other problems?

OpenStudy (anonymous):

Sure :) open a new thread :)

OpenStudy (anonymous):

I'll try anyway!

OpenStudy (anonymous):

Just @ me (@JoStudy) in your new question :) that way i'm sure to find you @nicole-0109